OpenStudy (anonymous):

A teacher gave her class two exams; 60% of the class passed the second exam, but only 48% of the class passed both exams. What percent of those who passed the second exam also passed the first exam?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

So this is exactly the inverse of the street light one. Remember how you set that one up?

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

A diner has collected data about customer coffee-drinking habits. They have calculated that P(cream) = 0.5, P(sugar) = 0.6, and P(cream or sugar) = 0.7. Determine the P(cream and sugar).

OpenStudy (anonymous):

correct ^_^

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

iddnt read that right

OpenStudy (anonymous):

\[ \textrm{P(a and b)} \ = \ P(a \cap b) \\ \textrm{P(a or b)} \ = \ P(a \cup b) \\ \ \\ P(a \cup b) = P(a) + P(b) - P(a \cap b) \]

OpenStudy (dan815):

yo explain why 80% of who passed 2nd exam passed the first one

OpenStudy (anonymous):

n * 6/10 = 48/100 n = 8/10

OpenStudy (anonymous):

oh, read that wrong

OpenStudy (anonymous):

yeah, 80% is totally not right. THat's just the percent of students who passed the first exam. oops.

OpenStudy (dan815):